OpenStudy (anonymous):

A certain region is losing about 11 million acres of rainforest each year. Find the linear function that calculates the changes in acres of rain forest in million of x years. Find f(5) and intercept the results

OpenStudy (ganpat):

let x be number of years.. Y = be total remaining area.. So Y = Y(x-1) -11(x)... what say ??

OpenStudy (ganpat):

for x = 5.. y = y (5-1) - 11(5) y = 5y - 5 - 55 4y = 60 y = 15 Million acers.. does that make sense ??

OpenStudy (anonymous):

Yes it does. I just didn't know where to plug in which number

OpenStudy (anonymous):

So over 5 years then they would lose 15 million acres?

OpenStudy (anonymous):

actually if they lose 11 million per year an after 5 years the 15 million does not make since.

OpenStudy (anonymous):

f(x) = x ....Where does the 11 fit in.

OpenStudy (anonymous):

so we use f(5). I know x equals 5 so on both sides x is replaced with the 5

OpenStudy (ganpat):

yes.. what i tried to figure out is, amount of ground remaining :-p.. thats also vague..

OpenStudy (ganpat):

simple w0uld be y = 11x, x = 5 so, y = 11 (5) = 55 acres,...

OpenStudy (anonymous):

I was think of something like f(5)=x-11...but I don't know what goes in front of the x to multiply by

OpenStudy (anonymous):

Okay so it would be represented by f(5)=11x

OpenStudy (ganpat):

no no.. x is number of months and it says 11acres per month...

OpenStudy (ganpat):

it was simple, i was trying to complicate.. sorry for that

OpenStudy (anonymous):

about 11 million acres per year

OpenStudy (anonymous):

per year would be represented by a 1

OpenStudy (anonymous):

and I am trying to find out with the use of f(x) over how many years and how many acres in that amount of time would disappear

OpenStudy (ganpat):

ya , the same and x is also number of years... sorry, TYPO error..

OpenStudy (anonymous):

Okay I think I have it. So it would be f(5)=11x f(5)=11(5) f(5)=55

OpenStudy (anonymous):

uggg...its to late to be doing this but it's do today.

OpenStudy (anonymous):

are you good with graphing.

OpenStudy (anonymous):

This next one has me boggled it gives me no pass through point.

OpenStudy (ganpat):

 check out this ..